### v2

# Package Bids in Combinatorial Electricity Auctions: Selection, Welfare Losses, and Alternatives

Abstract
A key challenge in combinatorial auctions is designing bid formats that accurately capture agents’ preferences while remaining computationally feasible. This is especially true for electricity auctions, where complex preferences complicate straightforward solutions. In this context, we examine the XOR package bid, the default choice in combinatorial auctions and adopted in European day-ahead and intraday auctions under the name “exclusive group of block bids”. Unlike parametric bid formats often employed in US power auctions, XOR package bids are technology-agnostic, making them particularly suitable for emerging demand-side participants. However, the challenge with package bids is that auctioneers must limit their number to maintain computational feasibility. As a result, agents are constrained in expressing their preferences, potentially lowering their surplus and reducing overall welfare. To address this issue, we propose decision support algorithms that optimize package bid selection, evaluate welfare losses resulting from bid limitations, and explore alternative bid formats. Our findings offer actionable insights for both auctioneers and bidders.
Keywords— Bidding/Auctions, Electricity Market, Stochastic Programming
1 Introduction
Electricity trading is constrained by complex technical requirements in transmission, production, and consumption, which underscores the need for auctions to facilitate efficient resource allocation and price discovery (Milgrom 2017). Most liberalized power markets feature at least a day-ahead auction, while some, such as the coupled European market, also offer multiple intraday auctions (Graf et al. 2024). In these auctions, the traded commodities are units of electric power (MW) for specified future periods. For example, a day-ahead auction might facilitate buying or selling electricity for the 8-9 am period tomorrow, while an intraday auction could allow trading for the 3:15-3:30 pm slot later within the day. This article addresses the following central question in electricity auction design: How can participants best express their preferences to buy or sell future electricity?
Preferences are expressed through bids, which can take various forms, but must conform to the bid formats defined by the auctioneer. Auctions that enable participants to bid on combinations of goods, rather than individual ones, are known as combinatorial auctions (Pekeč and Rothkopf 2003). Coupled European day-ahead and intraday auctions are examples of such combinatorial auctions, employing the default bid format for these types of auction, the XOR package bid, under the name “exclusive group of block bids” (Herrero et al. 2020).11endnote: 1Initially, block bids were limited to constant power profiles across periods, but this changed with the introduction of profile block bids (Herrero et al. 2020). Although terms for package and XOR bids vary between power exchanges, the concept remains consistent. Integrated into the EUPHEMIA market coupling algorithm, XOR bids are available for use by exchanges in all member states (NEMO Committee 2024b). However, they are most commonly utilized by EpexSpot and Nordpool, which operate in 16 European countries. In 2023, on average, 158 exclusive groups were submitted daily in the European coupled day-ahead auction (NEMO Committee 2024a). This format allows agents to bid on multiple power profiles that they are willing to buy or sell at a given price $p$, while ensuring that the auctioneer can accept at most one of these profiles (Figure 1(c)).
Figure 1: Illustration of an exclusive group of three block bids (XOR package bids).33endnote: 3The power exchange EpexSpot restricts block bids from including both sell and buy quantities for different time periods, which prevents arbitrageurs from utilizing XOR bids as demonstrated in Figure 1(b). This limitation is based on historical practices rather than any technical necessity. While so-called loop blocks enable the linkage of a buy and a sell profile (Karasavvidis et al. 2023), they cannot be grouped within an exclusive group. A better approach would be simply allowing a profile to contain both buy and sell quantities, eliminating this unnecessary restriction.
(a) Demand side.
(b) Arbitrageur.
(c) Supply side.
The rationale for using XOR package bids in electricity auctions stems from the intertemporal constraints that many participants face. For example, a battery storage system needs to charge before it can discharge, a shiftable load may prefer to consume in period x or y, but not both, and a thermal generator may need to operate over multiple periods continuously. Participants can ensure that these constraints are met by bidding on entire power profiles instead of separate quantities for each period. Declaring these package bids as XOR bids - grouping multiple profiles into an exclusive set - allows participants to avoid selling more power than they can produce or buying more than they can consume, as only one profile is accepted at most. XOR bids are capable of representing any type of preference (Nisan 2006), making them a versatile tool for agents on both the demand and supply sides, as well as for arbitrageurs (Figure 1(c)).
In combinatorial auctions, the auctioneer typically uses an optimization model to determine which bids to accept and the corresponding payments to agents. To ensure the computational tractability of this model, the auctioneer often limits the number of XOR package bids that can be submitted (Goetzendorff et al. 2015). This restriction creates a missing bids problem, where agents cannot fully express their preferences - some packages they are interested in cannot be bid on. As a result, agents can miss potential profits, leading to a loss in society’s welfare (Bichler et al. 2023). This issue is also evident in the coupled European day-ahead electricity auction, which imposes a limit of 24 block bids per exclusive group (EpexSpot 2024). In this work, we explore the effects of this limit and aim to answer the following questions:
  • How can participants select package bids that maximize their profit?44endnote: 4For simplicity, we use the term profit to encompass all types of participants. For demand-side agents, this typically refers to minimizing procurement costs.
  • What are the welfare implications of a limit on XOR bids?
  • Could alternative bid formats help solve the missing bid problem?
To address these questions, we must first examine the mechanisms employed by auctioneers to determine accepted bids and payments. Electricity auctions typically use a Walrasian mechanism (Milgrom and Watt 2024).55endnote: 5The widespread adoption of Walrasian mechanisms in electricity auctions is primarily driven by three factors: (i) the fairness and simplicity of uniform pricing, which ensures a single price per good for all agents (Bichler et al. 2018), (ii) its ability to serve as a clear scarcity signal, in contrast to nonlinear price signals that provide limited transparency (Ahunbay et al. 2024), and (iii) its facilitation of financial arbitrage across forward markets, from years ahead to real-time, through derivatives (Jha and Wolak 2023). Although specific implementations may vary, these mechanisms aim to find uniform prices that clear the market. That is,
  1. (i)
    Uniform prices (€/MW) are determined for electricity in each time period. Each agent pays the product of the accepted bid quantities and the uniform prices for each period.
  2. (ii)
    Only the most profitable XOR package bid is accepted; all are rejected if none are profitable.66endnote: 6For different bid formats, this rule may be stated differently. However, the main principle is that, at a Walrasian equilibrium, bids are accepted or rejected so that no agent would prefer a different outcome, given the equilibrium prices (Mas-Colell et al. 1995).
If the auctioneer consistently identifies such a Walrasian equilibrium and no agent has market power, which means that no agent can influence uniform prices, this mechanism is strategy-proof (Azevedo and Budish 2019). In a strategy-proof mechanism, truthful bidding becomes dominant, as agents need not strategize based on others’ bids. Furthermore, the allocation achieved under such conditions maximizes welfare in the auction (Mas-Colell et al. 1995).
However, the two ideal conditions - (i) a perfectly competitive market and (ii) the existence of a Walrasian equilibrium - rarely hold in practice. In the first case, this discrepancy requires the use of market power mitigation methods (Adelowo and Bohland 2024). For the second case, the Walrasian mechanisms employed can only approximate a Walrasian equilibrium, where not every agent is allocated their most profitable bid. While the European algorithm EUPHEMIA does not provide side payments to agents with such “paradoxically” outcomes (NEMO Committee 2024b), various forms of side payments are commonly incorporated into the mechanisms used by auctioneers in the United States (Stevens et al. 2024).
In addressing the above three questions on XOR bids, we assume those two ideals are true: (i) Walrasian equilibria exist, and (ii) no agent has market power. Although we acknowledge the severity of these assumptions, discussions with industry representatives involved in the European auction suggest that they are not uncommon. These assumptions are also made in Karasavvidis et al. (2024), seemingly the only study that addresses XOR bidding in electricity auctions.77endnote: 7Studies on XOR bid selection in other domains include Scheffel et al. (2012), which examines bounded rationality in spectrum auctions, or Hammami et al. (2021), which addresses the “bid construction problem” in transportation auctions. However, we did not find any direct transferable approach to electricity auctions. A key difference might be the high-quality price information available in daily electricity auctions. Furthermore, distinct auction mechanisms, such as pay-as-bid or multi-round formats, require specialized decision support strategies (Pekeč and Rothkopf 2003, Adomavicius et al. 2022). Combinatorial auctions are typically domain specific and often require customized solutions (Bichler et al. 2023). In their work, the authors formulate the XOR bid selection problem as a stochastic bilevel optimization model, where uniform prices are given as exogenous scenarios, and the lower level optimization reflects the bid acceptance rule under the assumed Walrasian equilibrium.88endnote: 8Karasavvidis et al. (2024) do not mention the possibility that this rule may not always hold, but instead implicitly assume equilibrium existence. Industry representatives note that paradoxically rejected or accepted bids are rarely an issue in the European day-ahead market. Although strong evidence for the frequent occurrence of Walrasian equilibria is lacking, some studies, such as a footnote in Reguant (2014) and a simulation by Meeus et al. (2009), suggest that this may be the case.
Given these two idealized assumptions, it seems logical to submit as many XOR bids as possible to hedge against uncertain prices. However, Karasavvidis et al. (2024) use their method to determine only ten package bids without mentioning the upper limit of 24. Similarly, an analysis of XOR bids in the German & Luxembourgian bidding zone in 2023 shows that many participants do not exploit this limit either (Figure 2(b)).99endnote: 9Perhaps unsurprisingly, a consultant from the energy industry shared that when they highlight the benefits of submitting the maximum number of bids, traders quickly see its value and wonder why they hadn’t thought of it sooner. The auction literature contains numerous examples where participants do not bid optimally due to a misunderstanding of auction mechanisms (Milgrom 2017). The concern that combinatorial auctions are too complex for agents to participate efficiently is as old as the idea of combinatorial auctions itself (Rassenti et al. 1982).
Figure 2: Usage of exclusive groups (XOR bids) in 2023 in the Germany & Luxembourg bidding zone.1111endnote: 11The historical bid data from the day-ahead auction was purchased from the power exchange EpexSpot.
(a) Demand bids.
(b) Supply bids.
It is crucial to highlight that day-ahead and intraday auctions are forward markets, requiring participants to account for opportunity costs and risk preferences in their bidding strategies. Opportunity costs may arise naturally within the market or from design flaws, as exemplified by INC-DEC gaming strategies (Graf et al. 2023). In this work, we do not address the estimation of opportunity costs or the handling of risk preferences. Instead, we assume risk-neutral participants who aim to maximize the expected profit based on a generic cost or valuation function1212endnote: 12A body of literature focuses on estimating opportunity costs, particularly in the context of hydropower; for further references, see Löhndorf and Wozabal (2023). Lastly, while extending our methods to convex risk measures, such as conditional value at risk, is conceptually straightforward, this is beyond the scope of the present work.
That said, the forward nature of electricity auctions requires participants to make rapid decisions as updated weather forecasts continually refine information about future intraday and real-time electricity prices.1313endnote: 13Moreover, there is a discussion about increasing the frequency of intraday auctions to enhance the overall efficiency of the European market, with a proposal to implement auctions every 15 minutes (Graf et al. 2024). This urgency demands that XOR bid selection methods be computationally efficient to avoid acting on outdated information. Despite this, Karasavvidis et al. (2024) formulated the bilevel model as a mixed-integer linear program (MILP), where binary variables grow linearly with the number of XOR bids and price scenarios. As a result, in one of their case studies, the model failed to terminate in an hour when determining only ten bids and considering just 20 price scenarios.
Our work introduces a linear programming (LP) approach for determining XOR bids. The polynomial time complexity of our method is crucial for two reasons: First, increasing uncertainty in electricity prices requires considering more price scenarios for robustness. Second, as we will argue later, the historical limit of 24 bids is likely outdated, with the number of allowable package bids expected to rise. We show that the stochastic bilevel model, maximizing expected profit under no market power and equilibrium assumptions, can be reformulated as an LP exploiting a total unimodular constraint matrix when the participant has a finite set of packages. If not, a finite subset of bids can be pre-selected by some heuristics.
When agents bid according to this algorithm, lack market power, and the auction accepts their most profitable bids, we can derive an upper bound on the profit loss caused by the limit on the number of bids. Moreover, assuming competitive equilibrium, we can extend this to bound the welfare loss for the entire auction. In particular, our bound is based on the Wasserstein distance between two probability measures: a discrete measure $\mathcal{Q}$ derived from the scenarios used for bid selection and a possibly continuous measure $\mathcal{P}$ representing the “true” price distribution. This relationship provides an interesting interpretation of the problem of bid selection and missing bids as a case of approximation of one probability measure with another (Figure 3(b)) - a well-explored topic in stochastic programming (Rujeerapaiboon et al. 2022).
Figure 3: An illustration of distributions $\mathcal{P}$ and their discrete approximations $\mathcal{Q}$.
(a) High uncertainty.
(b) Low uncertainty.
Notes: When uncertainty is high, $\mathcal{P}$ is more dispersed, requiring $\mathcal{Q}$ to use a larger number of scenarios for a good approximation. In contrast, when uncertainty is low, $\mathcal{P}$ is concentrated around a single mode, allowing $\mathcal{Q}$ to approximate it with fewer scenarios.
This theorem captures a key concept in microeconomic theory: In markets where equilibrium prices are known or can be easily determined through decentralized bargaining, no auctioneer is needed, as prices alone can guide allocations (Mas-Colell et al. 1995). In such cases, a single “bid” - indicating how much an agent is willing to produce or consume - is sufficient. However, when prices are unknown or difficult to determine, the auctioneer requires as much information as possible about agents’ preferences to guide efficient allocations (Milgrom 2017). Electricity markets lie somewhere in between these extremes. Although electricity prices are not fully known in advance, they usually follow predictable patterns (Lago et al. 2021). Therefore, although a single self-schedule based on forecasted prices might lead to inefficient or infeasible allocations, providing the auctioneer with detailed preferences might also be unnecessary if good price signals are available.
In general, equilibrium prices cannot be guaranteed in nonconvex markets, though near-equilibrium prices can often be approximated (see, e.g., Starr (1969) or Bichler et al. (2018)). However, unlike electricity auctions, agents in most other auctions typically lack prior knowledge of these prices. Consequently, studies addressing the missing bids problem - such as Goetzendorff et al. (2015) for TV ads and Bichler et al. (2023) for spectrum auctions - advocate for parametric bids, which work as follows: The agent submits parameters $\theta_{1},\ldots,\theta_{n}$, which the auctioneer uses to derive a set $\mathcal{M}(\theta_{1},\ldots,\theta_{n})$ of packages the agent is interested in, along with a function $f(\theta_{1},\ldots,\theta_{n}):\mathcal{M}(\theta_{1},\ldots,\theta_{n})\to\mathbb{R}$ that assigns bid prices to these packages, eliminating the need to enumerate all possible packages. These bids, commonly used for thermal plants and battery storage in US power systems, are known as “multi-part bids” and typically include parameters such as start-up costs or storage capacity (Herrero et al. 2020). A more direct approach would allow agents to submit sets $\mathcal{M}$ and functions $f$ directly, bypassing intermediate parameters. However, the auctioneer must restrict the complexity of $f$ and $\mathcal{M}$ to ensure tractable optimization, making this approach suitable only for agents with sufficiently simple preferences.1414endnote: 14The proposal by Bobo et al. (2021) suggests allowing agents to bid “state variables” and matrices defining a polyhedral set $\mathcal{M}$ and a convex piecewise-linear function $f$. This effectively allows any LP-representable $\mathcal{M}$ and $f$. To handle nonconvexities, the framework could extend to MILP-representable $\mathcal{M}$ and $f$.
The economics of most thermal power plants and battery storage systems can typically be captured through computationally feasible parametric or functional bids (Herrero et al. 2020). However, this may not apply to all resources, particularly emerging demand-side agents aggregating decentralized energy sources.1515endnote: 15Electrical energy is a complex commodity. Consequently, the systems that produce it, such as large thermal or hydropower plants, are intricate techno-economic systems that cannot always be easily described by a simple set of parameters or functional bid. On the demand side, human behavior further complicates matters, and aggregators such as utilities must account for the dynamics of their entire retail system when submitting buy bids in the wholesale market. The potential difficulty of parametric bids in accurately capturing preferences might be one reason why a large number of agents (about 40%, according to Byers and Eldridge (2023), for PJM in 2018) choose to self-schedule instead of using the available parametric bid formats. Another challenge is the willingness to participate. Complex bidding procedures often discourage participation, either due to a lack of understanding or a reluctance to share detailed operational data (Milgrom 2017). XOR bids offer a simpler alternative. They are easy to understand, require no operational data, and are applicable across all technologies. Their simplicity and versatility make them an ideal complement for agents for whom parametric and functional bids are unsuitable.
This article aims to help auction designers and agents using XOR bids understand and reduce the welfare loss caused by missing bids. To achieve this, Section 2 introduces our bid selection algorithm. Section 3 presents our Wasserstein theorem. In Section 4, we apply the algorithm to simulate the bidding behavior of a generator, battery storage, and a utility serving load in the German bidding zone during 2023. Section 5 demonstrates that package bids are not necessarily nonconvex, discusses aggregator and portfolio bidding, and briefly debates other types of package bids. Finally, Section 6 concludes the article.
2 Optimal Package Bid Selection
Consider an auction with $T$ periods, indexed by $t\in\mathcal{T}=\{1,\ldots,T\}$, and a participant with a valuation function $v:\mathbb{R}^{T}\to\mathbb{R}\cup\{-\infty\}$. The function $v(x)$ quantifies the agent’s cost $\big{(}v(x)<0\big{)}$ or utility $\big{(}v(x)>0\big{)}$ associated with the consumption or supply of power profile $x\in\mathbb{R}^{T}$. If $v(x)=-\infty$, the profile $x$ is infeasible for the agent, possibly due to technical constraints.
The agent may submit $B$ package bids in an exclusive group, where each bid is defined by $(x_{b},p_{b})$ for $b\in\mathcal{B}=\{1,\ldots,B\}$. Here, $x_{b}\in\mathbb{R}^{T}$ is a power profile of interest, and $p_{b}\in\mathbb{R}$ is the price the agent is willing to pay (or expects to receive) for $x_{b}$. The value $x_{bt}\in\mathbb{R}$ denotes the power consumed from the grid ($x_{bt}>0$) or supplied to the grid ($x_{bt}<0$) during period $t$, typically measured in MW. Bid prices $p_{b}$ are expressed in euros (€) in European day-ahead and intraday auctions.
2.1 Stochastic Bilevel Program
The agent determines XOR bids under the assumptions of no market power, no perfect information, and the expectation of an equilibrium outcome. This implies that the agent forecasts price scenarios $\lambda_{s}\in\mathbb{R}^{T},s\in\mathcal{S}=\{1,\ldots,S\}$ with probabilities $\pi_{s}$, and assumes the auctioneer selects the most profitable bid in each scenario:
$(x_{s}^{\ast},p_{s}^{\ast})\in\operatorname*{arg\,max}\big{\{}p_{b}-\langle\lambda_{s},x_{b}\rangle\ \big{|}\ (x_{b},p_{b})\in$
$\qquad\{(\mathbf{0},0),(x_{1},p_{1}),\ldots,(x_{B},p_{B})\}\big{\}}\quad\forall s\in\mathcal{S}.$ (1)
[…]
We assume a risk neutral agent aiming to determine $B$ package bids $(x_{b},p_{b}),b\in\mathcal{B}$, which maximize the expected profit. To do so, the agent seeks to solve:1616endnote: 16We assume the existence of an optimum, and therefore use $\max$ instead of $\sup$. However, it is important to note that this assumption does not necessarily hold, as no additional properties of $v$ have been specified.
$\max_{x_{b},p_{b}}$ $\quad\sum_{s\in\mathcal{S}}\pi_{s}\cdot\big{(}v(x_{s}^{\ast})-\langle\lambda_{s},x_{s}^{\ast}\rangle\big{)}$
[…]
This is a stochastic bilevel program, with the lower level defined by (2.1). It represents the “optimistic” version, where the upper level optimizes over the lower-level outcome $x_{s}^{\ast}$ (Kleinert et al. 2021). In reality, if there are multiple most profitable bids, the auctioneer, not the agent, makes the selection. The MILP formulation in Karasavvidis et al. (2024) can be derived from (2.1) using standard bilevel reformulation techniques (Kleinert et al. 2021). In the following, we show that this bilevel program can be reformulated as an LP.
2.2 Truthful Bidding
As discussed in the Introduction, a fundamental result in economic theory asserts that a mechanism leading to a Walrasian equilibrium guarantees truthful bidding as a dominant strategy, provided agents cannot influence the uniform price (Mas-Colell et al. 1995, Azevedo and Budish 2019). We show that this strategy-proofness extends to our setting, where agents cannot bid their valuation $v$ directly but must choose from a set of XOR package bids. To demonstrate this, we first analyze how program (2.1) would be affected by truthful bidding and provide proof that this is indeed a dominant strategy at the end of the section.
Assuming the agent bids truthfully, which means that they select a bid price equal to their valuation, that is, $p_{b}=v(x_{b})$ for all $b\in\mathcal{B}$, we can eliminate $p_{b}$ as a decision variable from problem (2.1). Consequently, the problem can be reformulated as:
$\max_{x_{b}\in\mathbb{R}^{T}}$ $\quad\sum_{s\in\mathcal{S}}\pi_{s}\cdot\big{(}v(x_{s}^{\ast})-\langle\lambda_{s},x_{s}^{\ast}\rangle\big{)}$
s.t. $\quad x_{s}^{\ast}\in\operatorname*{arg\,max}\big{\{}v(x_{b})-\langle\lambda_{s},x_{b}\rangle\ \big{|}$
$\qquad x_{b}\in\{\mathbf{0},x_{1},\ldots,x_{B}\}\big{\}}\quad\forall s\in\mathcal{S}.$ (3)
The objective function of the lower level now aligns with that of the upper level. As a result, the lower-level constraint can be omitted since the upper-level optimization will implicitly ensure the appropriate selection of $x_{s}^{\ast}$ (Kleinert et al. 2021):
$\max_{x_{b}\in\mathbb{R}^{T}}$ $\quad\sum_{s\in\mathcal{S}}\pi_{s}\cdot\big{(}v(x_{s}^{\ast})-\langle\lambda_{s},x_{s}^{\ast}\rangle\big{)}$
s.t. $\quad x_{s}^{\ast}\in\{\mathbf{0},x_{1},\ldots,x_{B}\}\quad\forall s\in\mathcal{S}.$ (4)
With these considerations in place, we can now establish that truthful bidding, i.e., $p_{b}=v(x_{b})$ for all $b\in\mathcal{B}$, is indeed a dominant strategy. As a consequence of this proposition, the agent can solve the single-level problem (2.2) instead of the bilevel problem (2.1) without any loss of expected profit.
Proposition 1.
Let $v^{1}$ and $v^{2}$ be the optimal values of program (2.1) without and with the constraints $p_{b}=v(x_{b})$ for all $b\in\mathcal{B}$, respectively. Then, $v^{2}\geq v^{1}$.
Proof.
Let $v^{3}$ be the optimal value of program (2.2). As shown above, $v^{3}=v^{2}$. Additionally, the feasible region for $x_{s}^{\ast}$ is larger in program (2.2) than in (2.1), where the lower-level constraint limits it. Thus, $v^{3}\geq v^{1}$, thereby concluding the proof. ∎
2.3 More Bids than Scenarios
Assume the agent is permitted to place more XOR package bids than the number of predicted price scenarios, that is, $B\geq S$. In this case, no optimization is necessary, as a solution to (2.2), and consequently to (2.1), can be obtained simply by executing the following algorithm:
(i) Forecast S price scenarios λs.(ii) For each scenario, determine an optimalprofile x¯sargmaxv(x)λs,x.(iii) Submit the package $\begin{array}[]{l}\text{(i) Forecast $S$ price scenarios $\lambda_{s}$.}\\ \text{(ii) For each scenario, determine an optimal}\\ \qquad\text{profile $\bar{x}_{s}\;\in\ \operatorname*{arg\,max}\;v(x)-\langle\lambda_{s},x\rangle.$}\\ \text{(iii) Submit the package bids}\\ \qquad\;\text{$\big{\{}\big{(}\bar{x}_{s},v(\bar{x}_{s})\big{)}\ \big{|}\ s=1,\ldots,S\big{\}}$.}\end{array}$ (10)
Our simulations in Section 4 show that using many price scenarios is beneficial in bid selection only when they are high quality and add meaningful information. For example, if only 20 out of 100 scenarios are reliable, focusing on these 20 is more effective than taking all 100 into account. In such cases, (2.1) can be efficiently solved using (10). Next, we consider another case where solving (2.1) is tractable.
2.4 Finitely Many Packages
Assume the agent is interested in only a finite number of packages. That is, there exists a finite set $\Sigma=\{\bar{x}_{1},\ldots,\bar{x}_{K}\}$ of profiles for which $v(x)\neq-\infty$, and $v(x)=-\infty$ for all $x\in\mathbb{R}^{T}\setminus\Sigma$. In this case, problem (2.2) is given by:
$\max_{x_{b}}$ $\quad\sum_{s\in\mathcal{S}}\pi_{s}\cdot\big{(}v(x_{s}^{\ast})-\langle\lambda_{s},x_{s}^{\ast}\rangle\big{)}$
s.t. $\quad x_{s}^{\ast}\in\big{\{}\mathbf{0},x_{1},\ldots,x_{B}\big{\}}\quad\forall s\in\mathcal{S}$
$\quad x_{b}\in\big{\{}\bar{x}_{1},\ldots,\bar{x}_{k},\ldots,\bar{x}_{K}\big{\}}\quad\forall b\in\mathcal{B}.$ (11)
The logic behind this program is as follows: First, the agent selects $B$ packages $x_{b}$ from the set $\{\bar{x}_{1},\ldots,\bar{x}_{K}\}$. Second, the agent chooses $S$ packages $x_{s}^{\ast}$ from the selected packages $\{\mathbf{0},x_{1},\ldots,x_{B}\}$ to represent the most profitable package in scenario $s$.
This logic can be captured using two types of binary variables. Let $\delta_{k}\in\{0,1\}$ denote whether the agent selects package $\bar{x}_{k}$ as one of its XOR bids, and let $\gamma_{ks}\in\{0,1\}$ denote whether package $\bar{x}_{k}$ is chosen as the most profitable package $x_{s}^{\ast}$. Importantly, $\gamma_{ks}$ can only be 1 if $\bar{x}_{k}$ is selected as one of the XOR bids, that is, $\gamma_{ks}\leq\delta_{k}$.
With these definitions, we can formulate (2.4) as the following binary program:
$\max_{\delta,\gamma}$ $\quad\sum_{s\in\mathcal{S}}\pi_{s}\cdot\big{(}v(x_{s}^{\ast})-\langle\lambda_{s},x_{s}^{\ast}\rangle\big{)}$ (12a)
s.t. $\quad x_{s}^{\ast}=\sum_{k\in\mathcal{K}}\gamma_{ks}\cdot\bar{x}_{k}\quad\forall s\in\mathcal{S}$ (12b)
$\quad\sum_{k\in\mathcal{K}}\gamma_{ks}\leq 1\quad\forall s\in\mathcal{S}$ (12c)
$\quad\gamma_{ks}\leq\delta_{k}\quad\forall s\in\mathcal{S},\>k\in\mathcal{K}$ (12d)
$\quad\sum_{k\in\mathcal{K}}\delta_{k}=B$ (12e)
$\quad\delta_{k},\gamma_{ks}\in\{0,1\}\quad\forall s\in\mathcal{S},\>k\in\mathcal{K}.$ (12f)
Here, $\mathcal{K}=\{1,\ldots,K\}$. Constraints (12c)–(12e) implement the selection logic of program (2.4). Specifically, (12e) ensures that the agent selects exactly $B$ packages to bid on, (12d) guarantees that only packages chosen by the agent can be selected as the most profitable, and (12c) ensures that at most one package is selected as the most profitable in each scenario.
To simplify this program further, we can eliminate $x_{s}^{\ast}$ by substituting (12b) into the objective function. Doing this, we obtain $v(\sum_{k\in\mathcal{K}}\gamma_{ks}\cdot\bar{x}_{k})$. Since at most one $\gamma_{ks}$ is 1 and the others are 0, we can draw the sum out of the function, resulting in $\sum_{k\in\mathcal{K}}\gamma_{ks}\cdot v(\bar{x}_{k})$. As $\bar{x}_{k}$ is a parameter of the optimization model, we can define $\bar{v}_{k}=v(\bar{x}_{k})$ as well as a parameter. Thus, we can formulate (12) as:
$\max_{\delta,\gamma}$ $\quad\sum_{s\in\mathcal{S}}\pi_{s}\cdot\sum_{k\in\mathcal{K}}\big{(}\bar{v}_{k}-\langle\lambda_{s},\bar{x}_{k}\rangle\big{)}\cdot\gamma_{ks}$
[…]
By examining the constraint matrix formed by constraints (12c)–(12f), it becomes clear that the binary restrictions are redundant. Therefore, the binary variables can be relaxed to continuous variables $\delta_{k},\gamma_{ks}\in[0,1]$ without sacrificing optimality.
Proposition 2.
Program (2.4) and its LP-relaxation are equivalent in the sense that their optimal values coincide, and there is a point $(\delta_{k}^{\ast},\gamma_{ks}^{\ast})_{s\in\mathcal{S},k\in\mathcal{K}}$ which is an optimal solution to both.
Proof.
It is straightforward to verify that the matrix formed by constraints (12c) - (12e) is totally unimodular by partitioning the rows (cf. Theorem 2.7 in part III.1 of Nemhauser and Wolsey 1988). Moreover, since adding identity matrices does not cancel total unimodularity, the polytope formed by (12c) - (12e) and $\delta_{k},\gamma_{ks}\in[0,1]\ \forall s\in\mathcal{S},\>k\in\mathcal{K}$ is integral (cf. Proposition 2.1 and Theorem 2.5 in part III.1 of Nemhauser and Wolsey 1988). Consequently, there is an integral vertex $(\delta_{k}^{\ast},\gamma_{ks}^{\ast})_{s\in\mathcal{S},k\in\mathcal{K}}$, which solves the LP relaxation and thus also (2.4). ∎
2.5 Infinitely Many Packages
We have identified two cases where (2.1) can be easily solved: when the number of bids allowed exceeds the number of scenarios, or when the agent is only interested in a finite set of packages. However, these conditions may not always hold, as electricity is a divisible good and thus agents could have an infinite number of packages they may be interested in.
The challenge in solving program (2.1) or its single-level equivalent (2.2) arises from the inclusion of $v(x)$, which can be an arbitrarily complex function. For example, it could represent the cost function of thermal power plants with intricate unit commitment dynamics or the utility function of an electric vehicle fleet operator shaped by its internal revenue management problem.
When $S>B$ and the agent has infinitely many packages of interest, we adopt a hybrid approach combining both strategies:
[…]
Although this does not guarantee an optimal solution to (2.1), it enables the agent to quickly determine package bids, even when $B$ and $S$ are large.
3 Bound on the Welfare Loss
Assume that no agent has market power and that there is an equilibrium price $\lambda$. If an agent could submit unlimited package bids, they could bid truthfully on every package:
$\big{\{}\big{(}x_{b},v(x_{b})\big{)}\in\mathbb{R}^{T+1}\ \big{|}\ v(x_{b})\neq-\infty\big{\}}.$ (20)
Given the equilibrium, the auctioneer would select the most profitable package as per equation (2.1), allowing the agent to achieve their maximum profit:1717endnote: 17Note that this is equivalent to the profit the agent could achieve with perfect knowledge of $\lambda$. As discussed in the introduction, an auctioneer is unnecessary if everyone knows the equilibrium price.
$\max_{x}\;v(x)-\langle\lambda,x\rangle.$
However, agents can only submit a subset of the bids in equation (20), which may prevent them from achieving their maximal profit. In this section, we develop a bound on the profit loss caused by limiting the number of bids.
Given our equilibrium assumption, the bound on agents’ surplus loss translates directly into a bound on the total welfare loss of the auction. This follows from the first welfare theorem in a partial equilibrium context, which states that when each agent maximizes their utility at the equilibrium price, total welfare is also maximized (Mas-Colell et al. 1995). Therefore, the loss in welfare at price $\lambda$ is the sum of the surplus loss of all agents.
3.1 Loss of Profit
The profit loss depends on the agent’s bidding strategy. We assume that the agent follows the simple strategy in (10) when fewer scenarios are considered than the allowed number of bids. This simplifies the theoretical analysis and serves as a lower bound compared to the more sophisticated strategies in (2.4) or (19). Given this strategy, the agent’s profit loss is:
$\Gamma(\lambda)\;=\;\max_{x}\;v(x)-\langle\lambda,x\rangle$ (21a)
$\qquad-\;\max_{x}\;\big{\{}v(x)-\langle\lambda,x\rangle\ \big{|}\ x\in\{\bar{x}_{1},\ldots,\bar{x}_{S}\}\big{\}},$ (21b)
where (21a) represents the agent’s profit if no bid restrictions exist, and (21b) is the profit achieved with strategy (10). The profit loss $\Gamma(\lambda)$ depends on the equilibrium price $\lambda$. If one of the $S$ price scenarios in strategy (10) is correct, then $\Gamma(\lambda^{\ast})=0$; otherwise, $\Gamma(\lambda)\geq 0$.
3.2 Ex Ante and Ex Post
Ex ante, the equilibrium price $\lambda$ is uncertain and can be described by a probability space $(\Omega,\mathcal{F},\mathcal{P})$, where $\Omega\subseteq\mathbb{R}^{T}$ is the sample space, $\mathcal{F}\subseteq 2^{\Omega}$ is a $\sigma$-algebra of events and $\mathcal{P}$ is a probability measure on $\mathcal{F}$. Ex post, the realized price follows a degenerate probability distribution, with $\mathcal{P}(\lambda)=1$ for a specific $\lambda\in\Omega$. The expected profit loss (ex-ante), as well as the actual profit loss (ex-post), are given by the expected value:
$\text{E}\big{[}\Gamma(\lambda)\big{]}=\int_{\Omega}\Gamma(\lambda)\ \mathcal{P}(d\lambda).$
The agent does not necessarily have prior knowledge of the probability distribution $\mathcal{P}$. Instead, following the bidding strategy described in (10), the agent considers $S$ distinct price scenarios $\lambda_{s}$, each of which occurs with a corresponding probability $\pi_{s}$. These scenarios collectively define a discrete probability distribution over the same measurable space $(\Omega,\mathcal{F})$:
$\mathcal{Q}(A)=\sum_{s\in\mathcal{S}}\pi_{s}\cdot\bm{1}_{A}(\lambda_{s})$
where $A\in\mathcal{F}$ is an event and $\bm{1}_{A}(\lambda_{s})$ the Dirac measure indicating whether $\lambda_{s}$ is in $A$ or not.
3.3 Wasserstein Bound
To bound the profit loss $\Gamma(\lambda)$, we establish a relation between the probability measures $\mathcal{P}$ and $\mathcal{Q}$. To achieve this, we employ probability metrics, which are commonly used to quantify the similarity between two probability distributions (cf. Rachev 2013). Among these metrics, the Wasserstein distance is particularly notable. Its popularity comes from its deep ties to the theory of optimal transport and its intuitive interpretation as the minimal cost of transporting mass from $\mathcal{P}$ to $\mathcal{Q}$ (cf. Villani 2009). Specifically, the Wasserstein distance of order 1 between $\mathcal{P}$ and $\mathcal{Q}$ is defined as:
$d_{W}(\mathcal{P},\mathcal{Q})=$ (22a)
$\ \inf_{\pi}$ $\;\int_{\Omega\times\Omega}\|\lambda_{1},\lambda_{2}\|_{2}\ \mu(d\lambda_{1},d\lambda_{2})$ (22b)
s.t. $\;\int_{\Omega}\mu(\lambda_{1},d\lambda_{2})=\mathcal{P}(\lambda_{1})\quad\forall\lambda_{1}\in\Omega$ (22c)
$\;\int_{\Omega}\mu(d\lambda_{1},\lambda_{2})=\mathcal{Q}(\lambda_{2})\quad\forall\lambda_{2}\in\Omega,$ (22d)
where $\|\cdot\|_{2}$ denotes the Euclidean norm. Using the Wasserstein distance $d_{W}(\mathcal{P},\mathcal{Q})$, we can give a bound on the profit loss as follows. The proof can be found in Appendix Appendix A.
Theorem 1.
Let $L=2\cdot\max\big{\{}\|x\|_{2}\ \big{|}\ x\in\mathbb{R}^{T},v(x)\neq-\infty\big{\}}$. Then $\text{E}\big{[}\Gamma(\lambda)\big{]}\leq L\cdot d_{W}(\mathcal{P},\mathcal{Q})$.
Theorem 1 highlights the interaction between price uncertainty and the number of bids $B$. When uncertainty is high, $\mathcal{P}$ is more dispersed, necessitating a larger number of bids $B$ for the agent to “obtain” a $\mathcal{Q}$ that closely approximates $\mathcal{P}$. Conversely, when uncertainty is low, $\mathcal{P}$ is more concentrated, allowing a good approximation even with fewer bids (Figure 3(b)).
4 Simulations: Generator, Storage, and Flexible Demand
So far, we have discussed an abstract auction with $T$ commodities and generic agents defined by their valuation $v(x)$. Now, we shift to explicit simulations using the 2023 German bidding zone in the coupled European day-ahead auction, where the commodities represent the 24 hours of the upcoming day. We use historical electricity prices from the ENTSO-E transparency platform (ENTSO-E 2024), assuming that they reflect competitive equilibrium prices, despite potential market power and paradoxically rejected bids. To illustrate that XOR bids work well for all types of participants, we consider three exemplary agents:
  • a thermal generator modeled with the classical unit commitment framework (Carrión and Arroyo 2006, Karasavvidis et al. 2021).
  • a battery storage system based on its standard model (De Vivero-Serrano et al. 2019, Karasavvidis et al. 2021).
  • a district heating utility employing heat pumps or gas boilers with demand-shifting capabilities through thermal storage, inspired by Bobo et al. (2021) and utilizing heat load data from Ruhnau et al. (2019).
All agents are modelled with an infinite number of potential packages of interest. Consequently, we apply algorithm (19) to select XOR package bids. We then analyze the profit loss of this bidding strategy, performing sensitivity analyzes on the number of bids, the number of scenarios, and the quality of price information.
The code to replicate these experiments can be accessed online at https://github.com/ThomasHubner/Bidding_on_Combinatorial_Electricity_Auctions. All experiments were run on a laptop with 32 GB RAM and an Intel i7-1260P processor. Optimization models were solved with Gurobi version 11 in standard settings. The execution time for Algorithm (19) was below 50 seconds for all cases considered.
4.1 Case Studies
Thermal generator
The plant can generate power within its minimum stable and maximum generation capacities and sell the output in the day-ahead auction. It can ramp up and down between hours, subject to a maximum ramping rate. Production costs consist of marginal costs, start-up and shutdown costs, and no-load costs. Additionally, when the plant is shut down (or started up), it must remain off (or on) for a predefined number of hours before it can start up (or shut down) again. A detailed model is provided in Appendix Appendix B, and the associated parameters are listed in Table 1(a).
Battery storage system
The battery can charge and discharge power, allowing it to both buy and sell in the day-ahead auction. Its state of charge must stay within its limits, and both charging/discharging efficiencies are considered. We assume that the battery’s state of charge at the end of the day must be the same as at the beginning of the day. The detailed model is given in Appendix Appendix C, with the parameters listed in Table 1(c).
District heating utility
We consider a utility that needs to supply a heat load using electric boilers, thus purchasing power in the day-ahead auction. The utility has aggregated heat storage available to shift the load between hours. It can also use a gas boiler if electricity prices are too high or reduce the load based on customer contracts, though it is not compensated for any curtailed load. The detailed model is provided in Appendix Appendix D, with the parameters listed in Table 1(b).
Table 1: Parameters used in the three case studies.
(a) Thermal generator
Description Value
No-load cost (€/h) 10,000
Marginal cost of 0-200 MW (€/MWh) 70
Marginal cost of 200-400 MW (€/MWh) 90
Marginal cost of 400-600 MW (€/MWh) 120
Start-up cost (€/start-up) 4,000
Shut-down cost (€/shut-down) 3,000
Minimum stable generation (MW) 100
Maximum power output (MW) 600
Ramp-up rate (MW/h) 200
Ramp-down rate (MW/h) 200
Minimum up-time (h) 4
Minimum downtime (h) 4
Initial operating state (MW) 0
Initial off hours (h) 0
(b) District heating utility
Description Value
Efficiency electric boiler 1
Efficiency gas boiler 0.9
Storage losses (%/h) 1
Cost gas (€/MWh) 90
Value of served load (€/MWh) 120
Capacity electric boiler (MW) 30
Capacity gas boiler (MW) 10
Capacity heat storage (MWh) 40
Maximum storage charging limit (MW) 20
Maximum storage discharging limit (MW) 20
Initial state-of-charge storage (MWh) 0
Hourly Heat Load (MW): 19, 20, 20, 21, 24,
32, 38, 36, 36, 35, 33, 32, 31, 31,
31, 32, 33, 33, 33, 33, 32, 29, 23, 20
(c) Battery storage system
Description Value
Maximum charging limit (MW) 10
Maximum discharging limit (MW) 10
Charging efficiency 0.9
Discharging efficiency 0.9
Minimum feasible state-of-charge (MWh) 0
Maximum feasible state-of-charge (MWh) 20
Initial state-of-charge (MWh) 10
4.2 Probabilistic Price Forecasts
The quality of bid selection depends on the accuracy of the forecasted price scenarios. We use a common two-step approach to generate these scenarios, widely adopted in both literature and industry (Löhndorf et al. 2013), and shown to perform competitively (Grothe et al. 2023). In the first step, we generate a point forecast for the price. In the second, we apply a post-processing method using the residuals from previous days to create the scenarios.
For the point forecast $y_{d}$ of the electricity price $\lambda_{d}$ on day $d$, we use an autoregressive model with LASSO regularization, implemented in the open-source Python package epftoolbox (Lago et al. 2021). This model is trained on German price data from 1/1/2019 to day $d$.
Using the point forecast $y_{d}$, we generate $S$ price scenarios $\lambda_{ds}$ with equal probability $\pi_{ds}=\frac{1}{S}$ by:
$\lambda_{d1}=y_{d},$ (23a)
$\lambda_{d2}=y_{d}-(y_{d-1}-\lambda_{d-1}),$ (23b)
$\lambda_{ds}=y_{d}-(y_{d-s-1}-\lambda_{d-s-1}),$ (23c)
where $(y_{d-1}-\lambda_{d-1})$ represents the forecast residual of the previous day. This approach generates $S$ price scenarios by adding the residuals of the last $S-1$ days to the point forecast $y_{d}$.
Since this method is simple to implement and publicly available, it serves as a lower bound on the price information accessible to bidders, who may use more advanced commercial forecasting methods.
4.3 Number of Bids
In the first set of experiments, we examine the profit loss as a function of the number of XOR package bids submitted by the agent. Using algorithm (19), we consider $S=150$ scenarios and bid numbers $B=1,2,5,10,20,50,100$. The case $B=1$ can be interpreted as a “self-schedule”, where the agent provides the auctioneer with only a single inflexible schedule.1818endnote: 18The term “self-schedule” is primarily used in US auctions to refer to an alternative to parametric bid formats, where asset owners, rather than the auctioneer’s algorithm, are responsible for dispatching the resource. Specifically, when parametric bids were introduced for battery storage systems, ISOs adopted a “self-schedule model” for these systems, allowing storage operators to directly manage dispatch. This approach was implemented in response to the potential inaccuracies of parametric bid formats for storage (Herrero et al. 2020). To ensure robust results, we randomly sample 100 days from 2023.
Figure 4 shows the average profit achieved with algorithm (19), expressed as a percentage of the maximum achievable profit (21a) based on the realized uniform price. Assuming equilibrium prices, the auctioneer accepts the most profitable bid as defined by (2.1). Missed profits in % are indirectly given by 100% minus the profit achieved presented in Figure 4.
Figure 4: Average profit by number of bids for 100 random days in 2023.
The results indicate that increasing the number of bids significantly improves profits. Generators and flexible loads, being less sensitive to price uncertainty, perform well with fewer bids. In contrast, battery storage systems, which rely on accurate price forecasts for effective temporal arbitrage, improve with higher bid counts and eventually catch up to the other participants.
4.4 Number of Scenarios
In the second set of experiments, we analyze the loss of profit as a function of the number of scenarios considered when selecting XOR bids by algorithm (19). Specifically, we generate $B=24$ package bids while accounting for $S$ = 24, 50, 100, 200, and 400 scenarios.
It is worth noting that for $S=24$, algorithm (19) simplifies to algorithm (10) due to $B\geq S$, resulting in package bids that correspond to a solution of the bilevel program (2.1).
To evaluate performance, we sample 100 days and compute the average profit, as detailed in Section 4.3. The results are summarized in Figure 5. While increasing the number of scenarios from 24 to 50 yields a significant improvement, further increases beyond 50 do not demonstrate a comparable benefit. This limitation can be attributed to our method used to generate scenarios from historical data (Section 4.2); as more scenarios are included, the model incorporates increasingly “outdated” information, diminishing its effectiveness.
Figure 5: Average achieved profit in relation to the number of scenarios.
4.5 Quality of Price Information
In the final set of experiments, we investigate the loss of profit as a function of the quality of the price forecasts. For this analysis, we fix $B=24$ and generate $S=100$ price scenarios $\lambda_{ds}$ for each day $d$, following the procedure in Section 4.2. To evaluate the impact of forecast accuracy, we progressively tighten the scenarios towards the actual price $\lambda_{d}$. Specifically, we define adjusted scenarios as $\lambda_{ds}^{\prime}=\lambda_{ds}-a\cdot(\lambda_{ds}-\lambda_{d})$, where $a$ is a scaling parameter.
This adjustment simulates improved forecast accuracy, as tightening scenarios closer to the true price can be interpreted as refining the forecasts. We consider five different values of $a$: $a=0,0.25,0.5,0.75,1$. The case $a=0$ corresponds to the baseline forecast used in previous analyses, while $a=1$ represents perfect information. To quantify the impact of these adjustments, we computed the Wasserstein distance between the discrete probability distribution formed by the 100 scenarios and the degenerate probability distribution centered on the actual price. This metric is plotted on the x-axis of Figure 6 for each value of $a$.
For the corresponding graph, we sampled 100 days and computed the average profit, as described in Section 4.3. The results show that as the accuracy of the forecast improves (i.e., as the Wasserstein distance decreases), profits increase, eventually converging to 100% in the case of perfect information.
Figure 6: Achieved profit in relation to the quality of price scenarios.
5 Discussions
If an auctioneer offers both parametric and XOR package bids, agents can choose the format that suits them. Agents whose preferences cannot be well represented by parametric bids or who prefer not to disclose detailed information can use the bid selection procedure in Section 2 to minimize profit loss. Agents whose preferences align with parametric bids can use them to completely avoid the risk of profit loss.
5.1 OR Package Bids
An alternative to parametric bids to address the missing bid problem in XOR bids is to use OR package bids (Bichler et al. 2023). Unlike XOR bids, which restrict acceptance to a single package, OR bids allow all submitted packages to be accepted independently. By submitting $B$ package bids, an agent effectively communicates interest in all $2^{B}$ combinations of these packages. However, OR bids are suitable only for agents with limited patterns of substitution in their preferences (Nisan 2006). This is because OR bids cannot express that two packages are substitutes, where accepting one excludes the other. This limitation is especially problematic in electricity auctions, where agents have preferences shaped by intertemporal constraints and upper limits on production or consumption. These constraints create strong substitution patterns: If too many bids are accepted, agents may exceed their capacity to deliver or consume, forcing costly after-auction adjustments. For agents such as shiftable loads or hydro reservoirs, substitution is even more critical - they operate based on producing or consuming in one period or another, but not both.1919endnote: 19The power exchange EpexSpot also permits “linking” multiple package bids, meaning a child bid is accepted only if its parent bid is accepted (Karasavvidis et al. 2021). However, linked package bids face the same limitation as OR bids, as they can only capture a narrow range of substitution patterns in the agent’s preferences.
Ironically, the EpexSpot power exchange allows agents to submit up to 100 OR package bids, but limits the XOR package bids to only 24 (EpexSpot 2024). Discussions with practitioners suggest that this restriction historically stems from the so-called flexi orders - a specific type of XOR bid. Flexi orders allowed agents to group several single commodity bids (consumption or production within a single time interval) into an exclusive set. For example, hydro reservoirs used this mechanism to let the auctioneer decide the optimal release time for stored water. Since a day consists of only 24 hours, grouping more than 24 hourly bids was unnecessary, leading to the 24-bid limit. However, when package bids are grouped, rather than single-commodity bids, this limit becomes obsolete. Computational concerns are unlikely the cause, given that up to 100 OR bids can be submitted.
Figure 7: Illustration of XOR bids with a minimum acceptance ratio of 0.
Notes: The upper profile is accepted by 50% whereas the lower profiles are accepted by 25%. In general, any convex combination of those profiles is possible if the minimum acceptance ratio of each profile is 0.
5.2 Convex Package Bids
Limiting the number of package bids is problematic because it is not causal; package bids only introduce a significant computational burden when they are nonconvex. However, since electricity is divisible, agents can specify a minimum acceptance ratio when submitting package bids. If this ratio is set to 0, the package bid becomes convex, as partial acceptance is allowed without restrictions. Formally, for each package bid $(x_{b},p_{b})$, the agent specifies a minimum acceptance ratio $\alpha\in[0,1]$. The actual acceptance rate $\delta_{b}$ of a bid must then satisfy $\delta_{b}\in\{0\}\cup[\alpha_{b},1]$. If multiple bids $b=1,\ldots,B$ are submitted as XOR bids, the constraint $\sum_{b\in\mathcal{B}}\delta_{b}\leq 1$ must be satisfied. Finally, the agent is allocated a package $x^{\prime}=\sum_{b\in\mathcal{B}}\delta_{b}\cdot x_{b}$, which may be a convex combination of its package bids (Figure 7).
Looking at the relative use of minimum acceptance ratios for package bids in the German bidding zone in 2023, we observe that 16.7% of the bids had a ratio of 0 (that is, convex bids), 2.7% were submitted with a ratio strictly between 0 and 1, and 80.6% of the bids had a ratio of 1 - a so-called “fill-or-kill” condition.9 While there are certainly agents with nonconvex preferences who need to submit package bids with $\alpha>0$, this clearly does not apply to every agent. Auctioneers could mitigate the welfare loss due to missing bids while still keeping optimization tractable by imposing a tight limit on package bids with $\alpha>0$, but allowing more package bids with $\alpha=0$.
5.3 Aggregators and Portfolios
Potential synergies between assets that consume or produce power encourage pooling them together. A notable example is the growing role of demand-side aggregators for distributed energy resources (Burger et al. 2017). To fully harness these synergies for societal welfare, it is necessary to allow coordinated bidding by such entities. XOR bids provide a practical format for this purpose, as designing parametric bids for aggregated resources is not a straightforward task (Liu et al. 2015, Herrero et al. 2020).
However, while synergies may justify bidding as a portfolio, concerns about potential market power abuse also arise. Ideally, large participants, such as major thermal power plants, would employ parametric bid formats to transparently disclose detailed cost information to the auctioneer, facilitating the detection of market power abuse (Adelowo and Bohland 2024). When participants avoid such transparency, for example, through self-scheduling or by opting for package or single-commodity bids, they impede the auctioneer’s ability to detect and mitigate market manipulation. Permitting large participants to cite “synergies” among assets as a rationale to avoid parametric bid formats thus risks reducing welfare (Ahlqvist et al. 2022).
6 Conclusions
Effective bid formats that allow agents to express preferences accurately are vital for efficient auctions. Auctioneers benefit from offering technology-specific parametric bid formats alongside universal and simple XOR package bids. To minimize welfare losses from limits on package bids, it helps to share as much information as possible about likely uniform prices. Allowing many convex package bids while placing tight limits only on nonconvex ones can keep optimization manageable while mitigating welfare losses.
Participants are encouraged to use parametric bid formats whenever possible. If these are not an option, it is advisable to submit the maximum number of packages allowed. When the number of desired bids exceeds the limit, bid selection algorithms can assist in prioritizing those with the highest expected profit.
This study assumes a competitive equilibrium and does not address price manipulation through market power (Reguant 2014) or exploitation of imperfect equilibrium outcomes (Byers and Eldridge 2023). Theoretical results by Starr (1969) and experiments such as those of Bichler et al. (2018) suggest that while uniform prices may not perfectly align with competitive equilibrium outcomes, they are generally also not too far off. Further research could help to better understand the welfare losses associated with imperfect bid formats when competitive equilibrium is not assumed.
Acknowledgements
The research published in this report was carried out with the support of the Swiss Federal Office of Energy (SFOE) as part of the SWEET PATHFNDR project. The authors bear sole responsibility for the conclusions and the results. The authors would like to thank Prof. Quentin Lété and Prof. Daniel Kuhn for their helpful comments and insightful discussions.
Appendix Appendix A Proof of Theorem 1
Denote the function in (21a) by $\psi(\lambda)$ and the function in (21b) by $\varphi(\lambda)$. The profit loss is then given by $\Gamma(\lambda)=\psi(\lambda)-\varphi(\lambda)$. In a first step, we show that both $\psi$ and $\varphi$ are Lipschitz-continuous with constant $L=\max\big{\{}\|x\|_{2}\ \big{|}\ x\in\mathbb{R}^{T},v(x)\neq-\infty\big{\}}$. In microeconomics, $\psi$ is known as the indirect utility function, which is known to be continuous if $v$ is continuous (Mas-Colell et al. 1995). However, we cannot necessarily assume $v$ to be continuous in electricity markets. Despite this, we can still establish the stronger property of Lipschitz continuity for $\psi(\lambda)$, as follows:
Lemma 1.
For all $\lambda_{1},\lambda_{2}\in\Omega$ holds $|\psi(\lambda_{1})-\psi(\lambda_{2})|\ \leq\ L\cdot\|\lambda_{1}-\lambda_{2}\|_{2}$.
Proof.
Let $\lambda_{1},\lambda_{2}\in\Omega$. Distinguish two cases: (i) $\psi(\lambda_{1})\geq\psi(\lambda_{2})$ and (ii) $\psi(\lambda_{1})<\psi(\lambda_{2})$.
Case (i): Let $x_{1}\in\operatorname*{arg\,max}v(x)-\langle\lambda_{1},x\rangle$. From $\psi(\lambda_{1})\geq\psi(\lambda_{2})$ and $\psi(\lambda_{2})\geq v(x_{1})-\langle\lambda_{2},x_{1}\rangle$ follows
$|\psi(\lambda_{1})-\psi(\lambda_{2})|$
$=\psi(\lambda_{1})-\psi(\lambda_{2})$
$\leq v(x_{1})-\langle\lambda_{1},x_{1}\rangle-v(x_{1})+\langle\lambda_{2},x_{1}\rangle$
$=\langle\lambda_{2}-\lambda_{1},x_{1}\rangle.$
Combining this with the Cauchy-Schwarz inequality, we arrive at
$|\psi(\lambda_{1})-\psi(\lambda_{2})|$ $\leq\ |\langle\lambda_{2}-\lambda_{1},x_{1}\rangle|$
$\leq\ \|x_{1}\|_{2}\cdot\|\lambda_{1}-\lambda_{2}\|_{2}.$
Case (ii): Let $x_{2}\in\operatorname*{arg\,max}v(x)-\langle\lambda_{2},x\rangle$. By similar arguments as in case (i) follows
$|\psi(\lambda_{1})-\psi(\lambda_{2})|\leq\|x_{2}\|_{2}\cdot\|\lambda_{1}-\lambda_{2}\|_{2}.$
Finally, the statement follows by $\forall\lambda\in\Omega$ and $\forall x\in\operatorname*{arg\,max}v(x)-\langle\lambda,x\rangle$ holds $\|x\|_{2}\leq\max\big{\{}\|x^{\prime}\|_{2}\ \big{|}\ x^{\prime}\in\mathbb{R}^{T},v(x^{\prime})\neq-\infty\big{\}}$. ∎
Lemma 2.
For all $\lambda_{1},\lambda_{2}\in\Omega$ holds $|\varphi(\lambda_{1})-\varphi(\lambda_{2})|\ \leq\ L\cdot\|\lambda_{1}-\lambda_{2}\|_{2}$.
Proof.
Follows by Lemma 1 and $\{\bar{x}_{1},\ldots,\bar{x}_{S}\}$ being a subset of every possible package. ∎
Lipschitz-continuity of $\psi$ and $\varphi$ is important as the Wasserstein distance is of so-called $\zeta$-structure which is closely linked to Lipschitz-continuity (Rachev 2013). With the help of Lemma 1 and Lemma 2, the proof of Theorem 1 can be carried out as follows:
Proof of Theorem 1 First of all, we relate $\text{E}\big{[}\Gamma(\lambda)\big{]}$ to distributions $\mathcal{P}$ and $\mathcal{Q}$ by:
$\text{E}\big{[}\Gamma(\lambda)\big{]}$ (25a)
$=\int_{\Omega}\Gamma(\lambda)\;\mathcal{P}(d\lambda)$ (25b)
$=\int_{\Omega}\Gamma(\lambda)\;\big{(}\mathcal{P}+\mathcal{Q}-\mathcal{Q}\big{)}(d\lambda)$ (25c)
Note that $\int_{\Omega}\Gamma(\lambda)\;\mathcal{Q}(d\lambda)=0$ since the bidding strategy (10) ensures that if a scenario $\lambda_{s}$ of the distribution $\mathcal{Q}$ occurs, then the agent is maximising its profit. Hence, we can continue by:
$=\int_{\Omega}\Gamma(\lambda)(\mathcal{P}-\mathcal{Q})(d\lambda)$ (25d)
$=\int_{\Omega}\Gamma(\lambda_{1})\;\mathcal{P}(d\lambda_{1})-\int_{\Omega}\Gamma(\lambda_{2})\;\mathcal{Q}(d\lambda_{2}).$ (25e)
By definition of $\mathcal{P}$ and $\mathcal{Q}$ as probability measures follows $\int_{\Omega}\mathcal{P}(d\lambda_{1})=\int_{\Omega}\mathcal{Q}(d\lambda_{2})=1$ and hence, we can continue (25e) by:
$=\int_{\Omega\times\Omega}\Gamma(\lambda_{1})\;\mathcal{P}(d\lambda_{1})\mathcal{Q}(d\lambda_{2})$
$-\int_{\Omega\times\Omega}\Gamma(\lambda_{2})\;\mathcal{P}(d\lambda_{1})\mathcal{Q}(d\lambda_{2}).$ (25f)
Since the Euclidean space is a Polish space, Theorem 4.1 from Villani (2009) ensures the existence of a solution $\mu^{\ast}(\lambda_{1},\lambda_{2})$ to (22). By definition, $\mu^{\ast}$ is a joint probability distribution with marginals $\mathcal{P}$ and $\mathcal{Q}$. Consequently, we can continue (25f) by:
$=\int_{\Omega\times\Omega}\Gamma(\lambda_{1})-\Gamma(\lambda_{2})\;\;\mu^{\ast}(d\lambda_{1},d\lambda_{2}).$ (25g)
Using the definition of the Lebesgue integral, this can be bounded above by:
$\leq\int_{\Omega\times\Omega}|\Gamma(\lambda_{1})-\Gamma(\lambda_{2})|\;\;\mu^{\ast}(d\lambda_{1},d\lambda_{2}).$ (25h)
By Lemma 1 and Lemma 2 follows that $\Gamma(\lambda)=\psi(\lambda)-\varphi(\lambda)$ is Lipschitz continuous with constant $L=2\cdot\max\big{\{}\|x\|_{2}\ \big{|}\ x\in\mathbb{R}^{T},v(x)\neq-\infty\big{\}}$. Hence, we can continue (25h) by:
$\leq L\cdot\int_{\Omega\times\Omega}\|\lambda_{1},\lambda_{2}\|_{2}\;\;\mu^{\ast}(d\lambda_{1},d\lambda_{2}).$ (25i)
Finally, by $\mu^{\ast}$ being a solution to (22) follows the statement.
Appendix Appendix B Thermal Generator
The thermal generator model used here is based on the classic unit commitment approach proposed by Carrión and Arroyo (2006), which was also applied by Karasavvidis et al. (2021) in the context of day-ahead auction participation of generators. The cost function $v(x)$ is defined by (26n), while the operational constraints governing the model are expressed in (26a)-(26m). For a detailed description of the model, see Carrión and Arroyo (2006) or Karasavvidis et al. (2021).
The power generated and sold by the thermal generator is the sum of the power output $g_{tk}$ from each block $k\in\mathcal{K}$ of the plant:
$x_{t}=\sum_{k\in\mathcal{K}}g_{tk}\qquad\forall t\in\mathcal{T}.$ (26a)
The production limit $P^{Max}_{k}$ for each block $k\in\mathcal{K}$ is enforced by:
$-g_{tk}\leq u_{t}\cdot P^{Max}_{k}\qquad\forall k\in\mathcal{K},\;\forall t\in\mathcal{T},$ (26b)
where $u_{t}$ is the binary variable describing whether the generator is committed or not. Remember that $x<0$ means selling power, leading to the minus in front of $g_{tk}$. The minimum stable generation $P^{Min}$ is enforced by:
$P^{Min}\cdot u_{t}\leq\sum_{k\in\mathcal{K}}-g_{tk}\qquad\forall t\in\mathcal{T}.$ (26c)
The ramp-up limit $R^{Up}$ and the ramp-down limit $R^{Dn}$ are enforced by:
$-R^{Dn}\leq x_{t-1}-x_{t}\leq R^{Up}\qquad\forall t\in\mathcal{T}.$ (26d)
Start-up costs $c^{up}_{t}$ and shut-down costs $c^{up}_{t}$ in scenario $s$ and period $t$ can be expressed by constraints:
$c^{up}_{t}\geq(u_{t}-u_{t-1})\cdot K^{Up}\qquad\forall t\in\mathcal{T}$ (26e)
$c^{dn}_{t}\geq(u_{t-1}-u_{t})\cdot K^{Dn}\qquad\forall t\in\mathcal{T},$ (26f)
where $K^{Up}$ represent the start-up and $K^{Dn}$ the shut-down cost. The initial-on-time $H^{Up}$ respectively initial-off-time $H^{Dn}$ is enforced by constraints:
$\sum_{t=1}^{H^{Up}}(1-u_{t})=0$ (26g)
$\sum_{t=1}^{H^{Dn}}u_{t}=0,$ (26h)
Finally, the minimum-up-time $T^{Up}$ and minimum-down time $T^{Dn}$ is enforced by the constraints:
$T^{Up}\cdot(u_{t}-u_{t-1})\leq\sum_{j=t}^{t+T^{Up}-1}u_{t}$
$\qquad\forall t\in[H^{Up}+1,\ldots,T-T^{Up}+1]$ (26i)
$-T^{Dn}\cdot(u_{t}-u_{t-1})\leq\sum_{j=t}^{t+T^{Dn}-1}(1-u_{t})$
$\qquad\forall t\in[H^{Dn}+1,\ldots,T-T^{Dn}+1]$ (26j)
$\sum_{j=t}^{T}u_{j}-(u_{t}-u_{t-1})\geq 0$
$\qquad\forall\;t\in[T-T^{Up}+2,\ldots,T]$ (26k)
$\sum_{j=t}^{T}(1-u_{sj})+(u_{t}-u_{t-1})\geq 0$
$\qquad\forall\;t\in[T-T^{Dn}+2,\ldots,T],$ (26l)
designed by Carrión and Arroyo (2006) to obtain a computationally efficient formulation. Finally, the constraints on the auxiliary variables are given by:
$g_{tk}\leq 0,\;u_{t}\in\{0,1\},\;c^{up}_{t},c^{dn}_{t}\geq 0$
$\qquad\forall k\in\mathcal{K},\;\;t\in\mathcal{T},$ (26m)
and the cost of power generation by:
$v(x)=\sum_{t\in\mathcal{T}}-K^{F}\cdot u_{t}-c^{up}_{t}-c^{dn}_{t}+\sum_{k\in\mathcal{K}}K^{M}_{k}\cdot g_{tk},$ (26n)
where $K^{F}$ are the no-load costs and $K^{M}_{k}$ are the marginal cost of production in block $k\in\mathcal{K}$.
Appendix Appendix C Battery Storage System
The following model for a battery storage system is commonly used for power systems applications of a battery, for instance, in De Vivero-Serrano et al. (2019) or Karasavvidis et al. (2023).
The buy and sell volume in hour $t$ and scenario $s$ is the difference between charging $g_{t}$ and discharging $d_{t}$:
$x_{t}=g_{t}-d_{t}\qquad\forall t\in\mathcal{T}$ (27a)
Charging and discharging are bounded by the charging respectively discharging limits $\overline{g}$ and $\overline{d}$:
$0\leq g_{t}\leq\overline{g}\cdot\delta_{t}\qquad\forall t\in\mathcal{T}$ (27b)
$0\leq d_{t}\leq\overline{d}\cdot(1-\delta_{t})\qquad\forall t\in\mathcal{T}$ (27c)
The binary variable $\delta_{t}\in\{0,1\}$ indicates whether the storage is charging or discharging. The state-of-charge $e_{t}$ is given by the balance equations:
$e_{t}=e_{t-1}+n^{g}\cdot g_{t}-\frac{d_{t}}{n^{d}}\qquad\forall t\in\mathcal{T},$ (27e)
where $n^{g}$ is the charging, and $n^{d}$ is the discharging efficiency. The state-of-charge must be kept between its lower $\underline{E}$ and upper $\overline{E}$ state-of-charge limit:
$-\underline{E}\leq e_{t}\leq\overline{E}\qquad\forall t\in\mathcal{T}$ (27f)
We assume that the state-of-charge must equal the initial $E_{0}$:
$e_{T}=E_{0}$ (27g)
Finally, we ignore degradation and maintenance costs, and the marginal costs are thus zero:
$v(x)=0$ (27h)
Appendix Appendix D District Heating Utility
The model of the district heating utility is oriented on the model in Bobo et al. (2021). The task of the utility is to serve the heat load $D_{t}$ in every hour $t$. For that, it can produce heat by a heat pump consuming power $x_{t}$, produce heat by a gas boiler consuming gas $y_{t}$, or curtail the load $z_{t}$. Moreover, it can discharge heat storage by $d_{t}$ or charge heat storage by $g_{t}$.
The heat balance equations are given by:
$\eta^{Hp}\cdot x_{t}+\eta^{Gas}\cdot y_{t}+d_{t}-g_{t}=D_{t}-z_{t}\qquad\forall t\in\mathcal{T},$ (28a)
where $\eta^{Hp}$ is the efficiency of the heat pump and $\eta^{Gas}$ is the efficiency of the gas boiler. The state-of-charge of the heat storage is given by:
$e_{t}=(1-\eta^{Loss})\cdot e_{t-1}+g_{t}-d_{t}\qquad\forall t\in\mathcal{T},$ (28b)
where $\eta^{Loss}$ is the percentage of heat energy lost between two hours. We assume that the state-of-charge must be, in the end, equal to the initial state-of-charge $E_{0}$:
$e_{Ts}=E_{0}$ (28d)
The state-of-charge, charge, and discharge, as well as the heat generated by the gas boiler and heat pump, are bounded
$e_{t}\in[0,\overline{e}],\;g_{t}\in[0,\overline{g}],\;d_{t}\in[0,\overline{d}],\;z_{t}\in[0,D_{t}],$
$\qquad x_{t}\in[0,\overline{x}],\;y_{t}\in[0,\overline{y}]\qquad\forall\;t\in\mathcal{T}$ (28e)
Finally, the valuation comprises the revenue from serving the load and the cost of gas:
$v(x)=K\cdot(D_{t}-z_{t})-C\cdot y_{t},$ (28f)
where $K$ is the value of the served load and $C$ is the cost of gas.